# Together We Go Further: LLMs and IDE Static Analysis for Extract Method Refactoring

### v1

@@ 4.4. RQ3 : Effectiveness of EM-Assist
EM-Assist- ($\mathcal{T}_{3}(e_{i})$) 60.6% 61.4% 66.4%
4.4.1. Subject systems and Experimental Setup
We selected six tools that are representative of a wide array of techniques, including four (JDeodorant (Tsantalis and Chatzigeorgiou, 2011), JExtract (Silva et al., 2015), SEMI (Charalampidou et al., 2017), and LiveRef (Fernandes et al., 2023)) that use static analysis-based rules and software quality metrics, and two (REMS (Cui et al., 2023) and GEMS (Xu et al., 2017)) that use prediction models based on machine learning techniques.
We first employ Community Corpus-A to evaluate the effectiveness of the selected tools. Following practices employed by the authors of other tools, we evaluate top-5 suggestions generated by the tools, calculating Recall@5 at tolerance levels of 1%, 2%, and 3%, (for details on evaluation refer to Section 4.3.1). For EM-Assist, we compute the metric for the best-performing hyper-parameters, including the ranking function $\mathcal{T}_{3}(e_{i})$, as identified in Section 4.3.1. For four tools (GEMS, JDeodorant, JExtract, SEMI), we reuse the evaluation results reported by other researchers (Xu et al., 2017) because they used the same dataset and they defined the same metric formulation as we do. For two tools (REMS and LiveRef), we had to run them ourselves: LiveRef had not previously been executed on Community Corpus-A, and REMS computes recall differently than any of the other tools. Notice that we contacted the authors and had extensive conversations with them to fully understand how to replicate their results best and how to use their tools so they perform the best.
Second, in order to bolster the robustness of our evaluation, we expand our evaluation to Extended Corpus, which replicates 2,849 Extract Method refactorings that took place in open-source projects. For this extended evaluation, we compare with LiveRef, the most recent tool, due to its ability to be run automatically on extensive corpora, setting it apart from other available tools, which require manual interventions and thus cannot be run at scale.
4.4.2. Results
Table 2 shows the comparative effectiveness of EM-Assist in relation to the other six tools. Compared to the other evaluated tools, GEMS has the second best results with Recall@5 of 54.2%, 59.8%, and 62.6% at tolerances from 1% to 3%, respectively. EM-Assist achieved a higher Recall@5 across all tolerance values ranging from 60.6% to 66.4%, which outperforms the second-best tool.
To further strengthen the validity of our results, we applied EM-Assist on the Extended Corpus that includes 2,849 actual refactorings from open-source projects. We applied the most recent tool, LiveRef, to the same dataset. For both tools, we calculated the recall for top-5 suggestions with a 3% tolerance value. LiveRef achieved recall of 6.5%, while EM-Assist’s recall was 42.1%, which is 6.4x times higher. These results show that EM-Assist’s effectiveness when applied to actual refactorings that developers perform in practice improves dramatically.
@@ 4.4.2. Results
When using a workbench of synthetic refactorings, EM-Assist has a superior recall rate compared to its peers. When using an oracle of real-world refactorings, EM-Assist’s recall rate improvements over its peers are 6.4x.
4.5. RQ4: Usefulness of EM-Assist
@@
EM-Assist is an interactive tool that provides maximum automation while still taking into account human input (see (Anonymous, 2023)). It presents refactoring suggestions to a developer, and based on their selection, EM-Assist applies the chosen refactorings and changes the code. Therefore, it is important to study the usefulness of these suggestions to developers and to gain insight into the decision-making process underlying their acceptance or rejection. We will take into account this feedback when releasing future versions of EM-Assist.
4.5.1. Data and Experimental setup
@@
To study the usefulness of the refactoring suggestions, we use the Firehouse survey research method (Murphy-Hill et al., 2015). A phenomenon and its solution are studied right after it happens (e.g., observing victims’ behavior right after a house catches fire). Accordingly, we engage with open-source developers immediately after they commit a long method into their repository, and we present them with suggestions for Extract Method. This immediate interaction allows developers to have fresh memories when responding to our questions, ensuring more reliable answers. Notably, this method, as demonstrated by Silva et al. Silva et al. (2016), boasts a significantly higher response rate compared to other survey-based studies that require developer participation. Even more important, unlike previous studies that asked third-party developers to assess the quality of refactoring suggestions on arbitrary pieces of code they are not familiar with, we instead go directly to the developers of the code. Because they are the most familiar with that recent code, they are the most qualified to judge the quality of the suggestion.
Figure 6. The firehouse method we used to evaluate EM-Assist.
@@
We evaluate the usefulness of EM-Assist’s suggestions by conducting firehouse surveys with developers who work on JetBrains/IntelliJ-Community Edition and JetBrains/Runtime, both are open-source projects. IntelliJ Community Edition is written in Java and Kotlin, has 15.8k stars, 417k commits, and 952 contributors, showcasing its maturity and making it an ideal candidate for our study. JetBrainsRuntime is written in Java, has 969 stars, 76k commits, and 806 contributors.
@@
Firehouse Survey Method: Figure 6 shows the process that we used for conducting firehouse surveys. We monitored the projects daily. We analyzed every new commit to identify newly added long methods. To enhance this process, we extended RefactoringMiner (Tsantalis et al., 2022) to prevent mistakenly identifying refactored methods (e.g., that are renamed or moved) as newly added. Once we identified a long method, we contacted the developer. To minimize the intrusion for the developers, we only asked their opinion for one single long method (regardless of how many methods they authored) and only presented three suggestions for their method (regardless of how many EM-Assist generated).
@@
After being presented with the top-3 refactoring suggestions generated by EM-Assist for their long method, we asked developers for their level of agreement with each suggestion (on a 6-point Likert scale ranging from Strongly Agree to Strongly Disagree). Additionally, we encouraged them to share their reasoning, detailing what aspects they liked or disliked about the suggestions or any changes they would consider.
Thematic Analysis: Since developers provide their motivations in free-form text, we employ thematic analysis, a widely used qualitative method for systematically extracting valuable information from their input. This approach, as recommended by Cruzes and Dybå Cruzes and Dybå (2011), helps identify and report patterns (themes) within the data, facilitating its deeper understanding. We followed the steps required by thematic analysis as suggested by other researchers (Braun and Clarke, 2006), in alignment with established research practices. Two authors of the paper independently read and assigned descriptive phrases (codes) to developers’ responses (Braun and Clarke, 2006). To ensure coding consistency between both labellers, they discussed and aligned their codes after labelling a random 25% sample of the data, as suggested by literature (Campbell et al., 2013). Subsequently, they coded the entire dataset of developers’ responses. After completing the process, the labellers finalized codes and derived thematic insights in line with established definitions of themes capturing data’s significant aspects.
4.5.2. Results
@@
We surveyed 20 developers, out of which 16 responded, bringing us to a 80% response rate. This response rate is notably higher than the typical response rate in questionnaire-based software engineering surveys, which typically hovers around 5% (Singer et al., 2008), and this can be attributed to us using firehouse surveys. Table 3 shows the results, specifically the agreement levels. For each developer survey, we present the entry for the highest-rated suggestion by the developer. Table 3 shows that even when giving just three ranked suggestions per method so that we do not overwhelm the developer, 81.3% of respondents find them useful (i.e., choosing a positive rating). This shows EM-Assist’s ability to generate useful refactoring opportunities that developers agree with. Since EM-Assist generates useful suggestions even for high-quality projects like the ones we used, we are confident it would discover useful refactoring opportunities in regular-quality projects.
@@
Table 3. Developers’ levels of agreement to the refactoring suggestions produced by EM-Assist
Strongly Agree Agree Somewhat Agree Somewhat Disagree Disagree Strongly Disagree
2 6 5 0 1 2
@@
Based on their qualitative comments, developers remarked that these refactoring suggestions significantly enhance code quality. They have also noted that EM-Assist’s suggestions provide fresh perspectives on their code, one developer stating “…these suggestions made me look at this code with new eyes once more, and I will try to refactor it in a different way”. Furthermore, developers express strong interest in seeing EM-Assist in production code, as exemplified by one developer’s gratitude: “Thank you for interesting suggestions! Hope to see this in production in the future.” These encouraging comments highlight the positive impact and potential usefulness of EM-Assist’s recommendations in the daily development process.
The top-3 reasons why developers did not accept EM-Assist’s suggestions are:
  1. (i)
@@
    The extracted method does not promote reuse: While this is a valid concern, it is important to note that EM-Assist focuses solely on individual methods as input and does not take into account broader contextual factors, such as a file or project-level considerations. In the future, we could extend EM-Assist’s scope to suggest extract methods for reuse at higher levels, such as a file, module, or project levels, for example, using techniques from (Alomar et al., 2023b, a).
A detailed analysis of these qualitative aspects can be found in our companion materials (Anonymous, 2023).
Developers confirm that EM-Assist’s suggestions are useful, with 81% agreeing with them.
5. Threats To Validity
  1. (1)
@@ 5. Threats To Validity
    Internal Validity: Does our work produce valid results? EM-Assist’s effectiveness can vary depending on the oracle used for evaluation. To address this, we use an independent, publicly available dataset that is widely used by researchers to assess similar tools. Our rigorous evaluation of the tool on this dataset demonstrates its superior performance in terms of recall when compared to other state-of-the-art tools. Moreover, we use two other datasets containing Extract Methods that open-source developers performed in reputable repositories. Furthermore, the firehouse survey results also show that open-source developers found EM-Assist’s suggestions useful.
  2. (2)
@@
    External Validity: Do our results generalize? EM-Assist’s effectiveness hinges on the chosen LLM model. As LLMs continue to advance, especially with upcoming iterations trained on more extensive datasets, we anticipate further improvements in outcomes. The architecture of EM-Assist allows for effortless adaptation to future LLMs. Lastly, it is important to highlight that while our approach is conceptually language-agnostic, our current EM-Assist implementation is tailored for Java and Kotlin. Therefore, we cannot extrapolate our performance results to programs written in other programming languages.
  3. (3)
    Verifiability: We make the EM-Assist, the datasets and results, publicly available (Anonymous, 2023).
6. Related Work
We group the related work into four categories: (i) static analysis and (ii) machine learning techniques for Extract Method refactoring, (iii) various studies on Extract Method refactoring, and (iv) tools that employ LLMs for various SE tasks.
@@ 6. Related Work
Static analysis: We present more details about the four static analysis tools for which we conduct a direct comparison in our evaluation (see Section 4.4). JDeodorant (Tsantalis and Chatzigeorgiou, 2011) automatically calculates block-based program slicing for variables in assignment statements. Given a slicing criterion as input, JDeodorant can extract relevant statements while preserving program behavior. JExtract (Silva et al., 2015) operates based on the block structure, where each block structure contains a group of statements organized with a linear control flow. Given a method, JExtract detects involved block structures and heuristically ranks them as refactoring candidates. SEMI (Charalampidou et al., 2017) explores the coherence between statements and returns code fragments with high cohesion as refactoring candidates. LiveRef (Fernandes et al., 2023) is an IntelliJ plugin that offers real-time Extract Method refactoring suggestions and immediate application, guided by code quality metrics and visual cues in the code editor. It continually adapts to code changes, providing live refactoring support.
Other researchers suggest Extract Method refactoring based on program slicing (Lakhotia and Deprez, 1998; Maruyama, 2001; Abadi et al., 2009). Others used software design principles, such as separation of concerns (Silva et al., 2014), or single responsibility principle (Charalampidou et al., 2017). Several others (Tairas and Gray, 2012; Yue et al., 2018; Yoshida et al., 2019; Alomar et al., 2023b) suggest Extract Method refactoring on the basis of eliminating code clones and code duplication. In contrast, we are using static analysis techniques to check and enhance the LLM results and to compensate for LLMs lacking program semantics for executing Extract Method correctly. Similar to (Charalampidou et al., 2017; Silva et al., 2015; Xu et al., 2017; Yang et al., 2009a; Fernandes et al., 2022) we employ a ranking mechanism for EM-Assist’s suggestions. Whereas previous work uses software quality metrics such as cyclomatic complexity, nesting depth and length, in our approach we rely on LLM-specific quality metrics, such as popularity and heat.
ML Models: We present more details about the two ML-based tools for which we conduct a direct comparison in our evaluation (see Section 4.4). REMS (Cui et al., 2023) utilizes multi-view representations from the code property graph. It then trains a machine learning classifier to guide the extraction of suitable lines of code as a new method. GEMS (Xu et al., 2017) encodes metrics related to complexity, cohesion, and coupling as features to train machine learning classifiers for recommending Extract Method refactoring opportunities.
@@
Other researchers worked on discovering the need for applying Extract Method refactoring. Aniche et al. (Aniche et al., 2020) evaluated the effectiveness of various ML algorithms for predicting software refactorings, including Extract Method. The models were trained on thousands of refactorings mined with RefactoringMiner (Tsantalis et al., 2022) from open-source projects. Van der Leij et al. (van der Leij et al., 2021) replicated the study (Aniche et al., 2020) at the ING company and discovered that ML models predict very well the opportunity for applying Extract Method refactoring. Others (AlOmar et al., 2022; Sagar et al., 2021) use commit messages to predict software refactorings. While these techniques are very good at predicting the type of refactoring a method needs to undergo, they complement nicely with EM-Assist and its ability to suggest code fragments to be extracted and also to apply the refactoring.
Miscellaneous Extract Method studies: Several studies (Negara et al., 2013; Murphy-Hill et al., 2012; Tsantalis et al., 2022) show that Extract Method refactoring is among the top-5 most frequently performed in practice, both for manual and automated refactoring. Murphy-Hill and Black (Murphy-Hill and Black, 2008) were some of the first researchers to uncover challenges that developers face when performing Extract Method refactorings.
LLMs in SE tasks: Hou et al. (Hou et al., 2023) conducted a comprehensive literature review on the application of LLMs in SE. White et al. (White et al., 2023) introduce patterns for prompting LLMs for various SE tasks, including code refactoring. In contrast, our work presents empirically-justified best practices for using LLMs for refactoring tasks.
7. Conclusions
@@
EM-Assist recommends Extract Method refactorings that align with developers’ preferences. This is evidenced when replicating thousands of real-world refactoring scenarios from open-source repositories. Moreover, our firehouse surveys with developers of high-quality codebases that authored recent changes revealed that 81.3% of respondents found EM-Assist’s suggestions useful.
@@
In this work, we focused on Extract Method refactoring for Java and Kotlin, but our approach can be expanded to other refactoring automation tasks, and to other programming languages. We hope that the best practices that we discovered for using LLMs effectively for refactoring tasks inspire others to further advance the field of refactoring.
8. Data Availability
We have made our tool and evaluation data publicly available on our website (Anonymous, 2023).

### v2

@@ IV-E2 Results
EM-Assist 57.6% $\pm$ 1.6 58.4% $\pm$ 1.6 63% $\pm$ 1.7
TABLE III: Evaluation results of EM-Assist with respect to six other tools on Extended Corpus of 1752 instances
Recall@5
Tool Tolerance 1% Tolerance 2% Tolerance 3%
EM-Assist $52.2\%\pm 0.9$ $52.7\%\pm 1$ 53.4% ±$\pm 1$
J-Extract 38.8% 39.3% 39.4%
@@
When using an oracle of synthetic refactorings, EM-Assist has a slightly superior recall rate compared to its peers. When using an oracle of real-world refactorings, EM-Assist’s recall rate improvements over its peers are 35%, showing it better aligns with developer preferences.
IV-F RQ4: Usefulness of EM-Assist
@@
EM-Assist is an interactive tool that provides maximum automation while still taking into account human input (see [42]). It presents refactoring suggestions to a developer, and based on their selection, EM-Assist applies the chosen refactorings and changes the code. Therefore, it is important to study the usefulness of these suggestions to developers and to gain insight into their decision-making process. We will take into account this feedback when releasing future versions.
IV-F1 Data and Experimental setup
@@
To study the usefulness of the refactoring suggestions, we use the Firehouse survey research method [30]. A phenomenon and its solution are studied right after it happens (e.g., observing victims’ behavior right after a house catches fire). Accordingly, we engage with open-source developers immediately after they commit a long method into their repository, and we present them with suggestions for EM generated by EM-Assist. This direct interaction with developers, leveraging their recent familiarity with the code, ensures they provide reliable answers and are best equipped to evaluate the quality of suggestions. Notably, this method, as demonstrated by Silva et al. [14], boasts a significantly higher response rate compared to other survey-based studies that require developer participation.
@@
To perform firehouse surveys, we engaged with developers contributing to the open-source projects JetBrains/IntelliJ-Community Edition [51] and JetBrains/Runtime [52]. The IntelliJ Community Edition, written in Java and Kotlin, has 15.8k stars, 417k commits, and 952 contributors. JetBrains/Runtime, written in Java, has 969 stars, 76k commits, and 806 contributors. These projects exemplify significant engagement, maturity, and attention to code quality, making them ideal for our study.
@@
Firehouse Survey Method: We monitored the projects daily, analyzing every new commit to identify newly added long methods. To enhance this, we extended RefactoringMiner [6] to prevent mistakenly identifying refactored methods (e.g., renamed or moved) as newly added. Once we identified a long method, we contacted the developer. To minimize the intrusion for the developers, we only asked their opinion for one single long method (regardless of how many methods they authored) and only presented three suggestions for their method (regardless of how many EM-Assist generated).
@@
Then we asked developers for their level of agreement with each suggestion on a 6-point Likert scale ranging from Strongly Agree to Strongly Disagree. Additionally, we encouraged them to share their reasoning, detailing what aspects they liked or disliked about the suggestions or any changes they would consider. As developers answered with free-form text, we conducted thematic analysis to interpret these responses, adhering to established best practices [53, 54, 55].
IV-F2 Results
@@
We surveyed 20 developers, out of which 16 responded, bringing us to a 80% response rate. This response rate is notably higher than the typical response rate in questionnaire-based software engineering surveys, which typically hovers around 5% [56], and this can be attributed to using firehouse surveys. Table IV shows the agreement levels. For each developer survey, we present the entry for the highest-rated suggestion by the developer. Table IV shows that even when giving just three ranked suggestions per method so that we do not overwhelm the developer, 81.3% of respondents find them useful (i.e., chose a positive rating). Thus, EM-Assist generates useful refactorings that align with developer preferences. As EM-Assist finds improvement opportunities and generates useful suggestions even for high-quality projects like the ones we used, we are confident it would discover useful refactoring opportunities in regular-quality projects.
@@
TABLE IV: Developers’ levels of agreement to the refactoring suggestions produced by EM-Assist
Strongly Disagree Somewhat Somewhat Agree Strongly
Disagree Disagree Agree Agree
2 1 0 5 6 2
@@
Developers remarked that these refactoring suggestions significantly enhance code quality. They appreciated a fresh perspective on their code: “…these suggestions made me look at this code with new eyes once more, and I will try to refactor it”. Furthermore, developers strongly desire to see EM-Assist in production: “Thank you for interesting suggestions! Hope to see this in production in the future.” These encouraging comments highlight the positive impact and usefulness of EM-Assist’s recommendations in the daily development process.
The top-3 reasons why developers did not accept suggestions:
  1. 1.
@@
    The extracted method does not promote reuse: While this is a valid concern, it is important to note that EM-Assist focuses solely on individual methods as input and does not take into account broader contextual factors, such as a file or project-level considerations. In the future, we could extend EM-Assist’s scope to suggest extract methods for reuse at higher levels, such as file, module, or project levels [57, 58].
Developers say 81% of EM-Assist’s suggestions are useful.
V Threats To Validity
  1. 1.
@@ V Threats To Validity
    Internal Validity: Does our work produce valid results? EM-Assist’s effectiveness can vary depending on the oracle used for evaluation. To mitigate this, we employ two datasets for assessment: a publicly available, independent dataset commonly utilized by other researchers, and a more extensive dataset derived from open-source repositories. Our thorough evaluation of EM-Assist on these datasets reveals its enhanced recall capabilities, outperforming other state-of-the-art tools. Furthermore, the firehouse survey results also show that open-source developers found EM-Assist’s suggestions useful.
  2. 2.
@@
    External Validity: Do our results generalize? EM-Assist’s effectiveness hinges on the chosen LLM model. As LLMs continue to advance, especially with upcoming iterations trained on more extensive datasets, we anticipate further improvements in outcomes. The architecture of EM-Assist allows for effortless adaptation to future LLMs. Lastly, it is important to highlight that while our approach is conceptually language-agnostic, our current EM-Assist implementation is tailored for Java and Kotlin. Therefore, we cannot extrapolate our performance results to programs written in other languages.
  3. 3.
    Verifiability: We make EM-Assist, the datasets and results, publicly available [31].
VI Related Work
We observed two primary approaches for EM refactoring: (i) tools designed to recommend specific code fragments for developers to refactor using EM, and (ii) tools that predict whether a host method needs EM refactoring. Accordingly, the related work is structured into two sections.
@@ VI Related Work
Suggesting EM refactoring: Several studies [4, 5, 6, 59] indicate that EM refactoring is among the top five most frequently performed practices, leading to the development of numerous supporting tools. JDeodorant [1] automatically calculates block-based program slicing for variables in assignment statements. Given a slicing criterion as input, JDeodorant can extract relevant statements while preserving program behavior. JExtract [27] operates based on the block structure, where each block structure contains a group of statements organized with a linear control flow. Given a method, JExtract detects involved block structures and heuristically ranks them as refactoring candidates. SEMI [8] explores the coherence between statements and returns code fragments with high cohesion as refactoring candidates. LiveRef [29] is an IntelliJ plugin that offers real-time EM refactoring suggestions and immediate application, guided by code quality metrics and visual cues in the code editor. It continually adapts to code changes, providing live refactoring support. REMS [12] utilizes multi-view representations from the code property graph. It then trains a ML classifier to guide the extraction of suitable lines of code as a new method. GEMS [11] encodes metrics related to complexity, cohesion, and coupling as features to train ML classifiers for recommending EM refactoring opportunities. Other tools propose EM refactoring based on program slicing [60, 7, 40], separation of concerns, and the single responsibility principle [8], while some [61, 62, 63, 57] aim to reduce code clones and duplication.
In contrast, our work pioneers the integration of LLMs with static analysis techniques to scrutinize and refine the outputs of LLMs, compensating for their lack of understanding in program semantics, thereby facilitating the correct and safe execution of EM refactoring.
@@
Predicting EM Refactoring: Researchers identified the necessity for implementing EM refactoring. Aniche et al. [64] evaluated the effectiveness of various ML algorithms for predicting software refactorings, including EM refactoring. Their models were trained on thousands of refactorings mined with RefactoringMiner [6] from open-source projects. Van der Leij et al. [65] replicated the study [64] at the ING company and discovered that ML models predict very well the opportunity for applying EM refactoring. Others [66, 67] use commit messages to predict software refactorings. While these tools are very good at predicting the type of refactoring a method needs to undergo, they complement nicely with EM-Assist and its ability to suggest code fragments to be extracted and also to apply the refactoring.
VII Conclusions
@@
EM-Assist recommends EM refactorings that align with developers’ preferences. This is evidenced when replicating thousands of real-world refactoring scenarios from open-source repositories. Moreover, our firehouse surveys with developers of high-quality codebases that authored recent changes revealed that 81.3% of respondents found EM-Assist’s suggestions useful.
We discovered a new set of best practices when using LLMs for refactoring. One of the biggest challenges with LLMs is taming non-determinism. Unlike in traditional static analysis tools where we avoid redundant computations, working effectively with LLMs requires re-prompting (asking the same question several times) and creating a superset of all suggestions. This was key to get the most out of the LLM. Thus, we designed novel ranking methods to take into account the LLM workflow. Moreover, we learned that the more precise the prompt, the higher the quality of the suggestions. Few-shot learning worked best for refactoring tasks.
@@
In this work, we focused on EM refactoring for Java and Kotlin, but our approach can be expanded to other refactoring automation tasks, and to other programming languages. We hope that the best practices that we discovered for using LLMs effectively for refactoring tasks inspire others to further advance the field of refactoring.
VIII Open Science Policy
We have made EM-Assist and evaluation data available anonymously on our website [31], and the tool is freely available to be installed from the JetBrains Marketplace [42].
IX Acknowledgements
We thank the ML Methods in Software Engineering Lab at JetBrains Research, and the FSE-2024 reviewers for their insightful and constructive feedback for improving the paper. This research was partially funded through the NSF grants CNS-1941898, CNS-2213763, and the Industry-University Cooperative Research Center on Pervasive Personalized Intelligence.